OpenStudy (anonymous):

Find the derivative of y=cos(2/x). I know the answer I just need to know the process.

geerky42 (geerky42):

Chain rule: \(\dfrac{d}{dx}f(g(x)) = f'(g(x))g'(x)\) So \(\dfrac{d}{dx}\cos(2/x) = -\sin(2/x)\cdot-2/x^2 = \boxed{\dfrac{2\sin(2/x)}{x^2}}\)

OpenStudy (anonymous):

How did you find the derivative of \[\frac{ 2 }{ x }\]

geerky42 (geerky42):

\(\dfrac{2}{x}=2x^{-1}\) Use power rule here.

OpenStudy (anonymous):

Thank you. Sometimes I overthink things and it really effects my understanding of calculus.

geerky42 (geerky42):

No problem